Linear and nonlinear outgrowths of the inverse square potential

Jerome Goldstein Univ. of Memphis, USA

Linear outgrowths include new L^{p} versions of Hardy’s inequality for parametric extensions of the Schroedinger operator with the inverse square potential, and their application to semigroup generation. Nonlinear outgrowths and applications include nonexistence of positive solutions for a variety of parabolic problems.